# DAViD: Modeling Dynamic Affordance of 3D Objects using Pre-trained Video Diffusion Models

### v1

@@ Abstract
Understanding the ability of humans to use objects is crucial for AI to improve daily life. Existing studies for learning such ability focus on human-object patterns (e.g., contact, spatial relation, orientation) in static situations, and learning Human-Object Interaction (HOI) patterns over time (i.e., movement of human and object) is relatively less explored. In this paper, we introduce a novel type of affordance named Dynamic Affordance. For a given input 3D object mesh, we learn dynamic affordance which models the distribution of both (1) human motion and (2) human-guided object pose during interactions. As a core idea, we present a method to learn the 3D dynamic affordance from synthetically generated 2D videos, leveraging a pre-trained video diffusion model. Specifically, we propose a pipeline that first generates 2D HOI videos from the 3D object and then lifts them into 3D to generate 4D HOI samples. Once we generate diverse 4D HOI samples on various target objects, we train our DAViD, where we present a method based on the Low-Rank Adaptation (LoRA) module for pre-trained human motion diffusion model (MDM) and an object pose diffusion model with human pose guidance. Our motion diffusion model is extended for multi-object interactions, demonstrating the advantage of our pipeline with LoRA for combining the concepts of object usage. Through extensive experiments, we demonstrate our DAViD outperforms the baselines in generating human motion with HOIs.
@@
Figure 1: Dynamic Affordance. Given an input 3D object (left), our method learns Dynamic Affordance to model the dynamic patterns that occur during HOI (right). To build it, we present a framework to synthesize diverse 4D HOI samples (middle) by leveraging a pre-trained video diffusion model.
@@ 1 Introduction
Humans effortlessly know how to appropriately utilize objects to achieve their goals in a given environment. This knowledge, also known as affordance [22], encapsulates a comprehensive understanding of how humans interact with specific objects. This affordance should consider not only the patterns such as contact, orientation, and position appearing in Human-Object Interaction (HOI) at a specific time instance but also the dynamic movements of humans and objects during interactions. For example, even when positioned directly behind the shopping cart, holding its handle and facing forward, we push the shopping cart ahead instead of lifting it up. Learning such affordances informs agents about how humans use objects and helps them act proactively within their environments, making it crucial in the fields of robotics and AI. Given its importance, numerous studies on learning affordances [93, 84, 79, 92, 57, 72, 42, 3, 63, 61, 64, 78, 65, 16, 12, 120, 11, 112, 122, 97, 52, 62, 71, 121, 14, 75, 38, 55, 13] appeared.
However, previous studies on affordance learning focus on patterns that occur in static HOIs. Some studies infer affordances in a given static context (typically provided as images) through contact on 2D images [48, 4], 3D humans [90, 109], and 3D objects [108, 109]. Others [27, 39] represent affordances as a distribution by aggregating patterns from various generated static HOI samples. While recent studies generate human motions including HOI [49, 106, 70, 50, 107, 104] and Human-Scene Interaction (HSI) [100, 111, 104], they focus on limited human motions (e.g., stand and sit or pick and place) regardless of the object categories, which lack in capturing dynamic patterns appearing during real-world HOIs, which differ significantly depending on the object category.
In this paper, we introduce Dynamic Affordance, a novel type of affordance that models dynamic patterns in HOIs. The motivation for this concept arises from the observation that there are not only static patterns but also tendencies in the movements of both humans and objects over time during HOI. For instance, as shown in Fig. 2, we typically push a shopping cart and pull a suitcase while walking forward. We rarely lift the handles of the shopping cart and suitcase vertically. However, traditional studies based on static affordance (e.g., relative position, contact) do not fully capture how humans interact with objects over time. We address this by training our DAViD which models the distribution of dynamic movements for both a human and an object during HOIs. Importantly, as we focus on modeling HOI motion depending on the object category, we treat human motion interacting with a specific object as a single concept and generate human motion faithful to that concept. To achieve this, we extend the human motion space by attaching Low-Rank Adaptation (LoRA) to a pre-trained motion diffusion model (MDM). Then, by modeling the remaining object motion space conditioned by human motion using a score-based diffusion model, we finally construct our Dynamic Affordance.
To train our DAViD, we need a dynamic human-object pair dataset involving diverse interactions. While most existing 3D HOI datasets provide dynamic sequences, they only focus on specific object categories [98, 35, 115] (i.e., chair, ball), captured by marker-based capture methods [29, 118, 86, 19] which are difficult to scale. Furthermore, most existing datasets are captured in lab-controlled environments [5, 34, 40], making it challenging to obtain HOI data for various object categories. Our key observation for this is that a pre-trained video generation model can represent dynamic patterns of object usage in the form of videos. As shown in Fig. 2, the pre-trained video generation models already know how to generate videos with plausible dynamic patterns (e.g., pushing a shopping cart, pulling a suitcase). However, leveraging a pre-trained video generation model to learn Dynamic Affordance presents two challenges: (1) the generated video should include HOI with an object that is geometrically identical to the given 3D object, and (2) both human and object motions in the video should be uplifted into a common 3D space. To address these challenges, we propose a Canny Edge [8] guided HOI image generation combined with an uplifting pipeline based on 2D correspondences.
We validate the efficacy of our approach by learning Dynamic Affordance on 30 3D object meshes from diverse categories, gathered from multiple sources [5, 34, 10, 105, 60, 39, 2]. We compare our Dynamic Affordance against existing methods [49] on the FullBodyManipulation dataset [49], showing that our model using synthetic samples, outperforms competitors on generating human motion including HOI. Additionally, we verify the advantage of Dynamic Affordance on generating human motion with multi-object interaction by training diffusion models using Dynamic Affordance. We will release the results and our source code.
In summary, our main contributions are summarized as: (1) we introduce a new type of affordance named Dynamic Affordance, which captures the distribution of both human motion and human conditioned object pose during interactions; (2) we propose a pipeline to generate diverse 4D HOI samples from 3D object mesh without laborious capturing; (3) we present a method for modeling the dynamic movements of humans in HOI as a single concept using LoRA and highlight its advantages in generating HOI motions that interact with multiple objects.
@@
Figure 2: Dynamic Patterns in HOI. People show dynamic patterns during HOI. Pre-trained video diffusion model has the knowledge of the dynamic patterns.
@@ 2 Related Work
Reasoning Visual Affordance. Affordance is a concept defined by Gibson [22] as the set of possible interactions that an agent can perform within an environment. As the concept narrows to the primary agent (or humans) and the specific interaction target (or objects), the area of learning affordance mainly focuses on HOI. Specifically, followed by the intuition that the visual geometry of an object is closely related to its functionality, many studies leverage visual cues to reason about object’s affordance. The studies are subdivided into various areas, including affordance classification [93, 84, 79, 92, 57, 72, 42], affordance detection [3, 63, 61, 64, 78, 65, 16, 12], and affordance segmentation [75, 38, 55, 13]. Such traditional studies on affordance infer affordance for the given context (mostly provided as images), as a form of action labels [7, 32, 46], and contact regions in 2D [48, 4] or 3D [90, 108]. Later studies model static HOI patterns (e.g., contact) during interactions [15, 27, 39], going beyond the understanding of individual HOI situations. Recent studies synthesize human motion including HOI [106, 70, 100, 111, 50, 107, 104], showing potential to learn dynamic HOI patterns, but they focus on a text-driven approach, generating quite consistent human motions regardless of object categories (e.g., dragging a chair, holding a backpack). In contrast, our method focuses on learning dynamic HOI patterns of objects, observable in real-world HOIs.
@@
Learning Concepts in Diffusion Models. Concept learning, also known as concept customization, is a strategy to extend the space of the pre-trained model to understand specific concepts (e.g., subjects, styles). Some studies tune the embeddings [21, 95] corresponding to a given concept in a pre-trained diffusion model, while others [76, 44] tune both the model weights and embeddings together. Recently, LoRA [33], originally used in large language models (LLMs), shows impressive results in producing images well aligned to the given concept [77], leading to an increase of attempts [80, 26, 82] to use LoRA for concept learning. Such concepts can be provided in various forms, including images [51, 119, 53, 24], videos [102, 26], and audios [96]. However, concept learning in fields beyond images, video, and audio is relatively less explored. We apply LoRA to the human motion space of pre-trained MDM [88] to learn the dynamic patterns displayed in HOI as a single concept, extending the original human motion space in terms of HOI.
@@ 3 Method
Our approach aims to learn the dynamic affordance of a target object from synthetically generated 2D videos via a video diffusion model [1]. In the first stage, we generate diverse 4D HOI samples given a 3D object mesh of the target (Sec. 3.1). Then, our method models the collected 4D HOI samples into our DAViD, which consists of (1) a motion diffusion model to synthesize feasible human motions related to the target object along with LoRA, and (2) an object pose diffusion model to synthesize dynamic object motions paired with the synthesized human motion (Sec 3.2). An overview of our method is shown in Fig. 3.
@@
Our pipeline begins with a 3D mesh input for the target object and produces diverse 2D HOI videos showing how a person interacts with (or uses) the target object.
2D HOI Video Synthesis. Given an input 3D object mesh, we render it from multiple views using cameras $\{\Pi_{c}\}_{c=1}^{C}$. For the intrinsic parameters, we follow the perspective camera model of the world-grounded human mesh recovery method [81] which is used in our 3D lifting process, so that we can have a consistent camera setup. For the extrinsic parameters (camera viewpoints) $\{\bm{R}_{c},\bm{t}_{c}\}_{c=1}^{C}$, we consider diverse commonly observable viewpoints for each object type (either ground-placed object or portable) with default object pose $\{\bm{R}_{o},\bm{t}_{o}\}$, similar to the previous approach, ComA [39]. Specifically, for the relatively stationary ground-placed object (e.g., motorcycle), we consider horizontal viewpoint variations centered on the target object, while for the portable objects (e.g., trumpet), we render from the sampled viewpoint within a range. The object type is automatically inferred via a vision-language model [67], by querying with the rendered image with a query prompt, “Is the object plausible to be on the ground while it is being used?”.
After rendering the object, we synthesize a realistic human that naturally interacts with the rendered 2D object via an off-the-shelf pre-trained 2D diffusion model [45]. Unlike the previous approaches [43, 110, 39] for a similar goal that uses inpainting by providing a mask region where the inpainting model to fill, we synthesize the entire image via ControlNet [114], conditioning on the edge cues of rendered object [8]. Our primary motivation for this strategy is to synthesize not only plausible human-object interaction for the given object viewpoint but also the realistic background scene, which is important for 3D lifting, while preserving the pose and shape of the rendered object.
Once we synthesize HOI images for the target object from many views, we refine the generated 2D HOI images by manually filtering out low-quality or undesirable ones. As we only need a few-shot sample set to train our LoRA, this process requires minimal human effort, taking approximately three minutes per category.
Finally, we synthesize corresponding 2D HOI videos ${\{\mathcal{V}_{d}\}_{d=1}^{N}}$ from the 2D HOI images ${\{\mathcal{I}_{d}\}_{d=1}^{N}}$ by using a pre-trained image-to-video generation model [1] with the same text prompt used for generating images. It is noted that our synthesized 2D HOI videos have advantages over in-the-wild 2D real videos, which is crucial for 4D HOI sample generation: (1) our HOI videos contain diverse views and HOI scenarios for the known target 3D object, and (2) we know the corresponding camera parameters and the 3D object pose of the initial frame of the video.
@@
Uplifting for 4D HOI Sample Generations. To uplift human motions into 3D from the synthesized 2D videos, we use an off-the-shelf world-grounded human motion recovery models, GVHMR [81] and TRAM [99], denoted as $\mathbf{F}_{\text{human}}$. We use GVHMR as the default method for module $\mathbf{F}_{\text{human}}$ except in the scenarios where human translation occurs without walking motions (e.g., riding a motorcycle). In such cases, we use TRAM [99] instead, as GVHMR [81] typically fails to predict plausible camera motion under these conditions. Therefore,
$\mathbf{F}_{\text{human}}(\mathcal{V}_{d})=\{\bm{\theta}_{d}^{f_{i}},\bm{\beta}_{d}^{f_{i}},\bm{R}_{d}^{f_{i}},\bm{t}_{d}^{f_{i}},\bm{R}_{\text{cam},d}^{f_{i}},\bm{t}_{\text{cam},d}^{f_{i}}\}_{i=1}^{F},$ (1)

### v2

@@ Abstract
Modeling how humans interact with objects is crucial for AI to effectively assist or mimic human behaviors. Existing studies for learning such ability primarily focus on static human-object interaction (HOI) patterns, such as contact and spatial relationships, while dynamic HOI patterns, capturing the movement of humans and objects over time, remain relatively underexplored. In this paper, we present a novel framework for learning Dynamic Affordance across various target object categories. To address the scarcity of 4D HOI datasets, our method learns the 3D dynamic affordance from synthetically generated 4D HOI samples. Specifically, we propose a pipeline that first generates 2D HOI videos from a given 3D target object using a pre-trained video diffusion model, then lifts them into 3D to generate 4D HOI samples. Leveraging these synthesized 4D HOI samples, we train DAViD, our generative 4D human-object interaction model, which is composed of two key components: (1) a human motion diffusion model (MDM) with Low-Rank Adaptation (LoRA) module to fine-tune a pre-trained MDM to learn the HOI motion concepts from limited HOI motion samples, (2) a motion diffusion model for 4D object poses conditioned by produced human interaction motions. Interestingly, DAViD can integrate newly learned HOI motion concepts with pre-trained human motions to create novel HOI motions, even for multiple HOI motion concepts, demonstrating the advantage of our pipeline with LoRA in integrating dynamic HOI concepts. Through extensive experiments, we demonstrate that DAViD outperforms baselines in synthesizing HOI motion.
@@
Figure 1: DAViD for Learning Dynamic Affordance. Given an input 3D object (left), our method, DAViD, learns Dynamic Affordance to model the dynamic patterns that occur during HOI (right). To build it, we present a framework to synthesize diverse 4D HOI samples (middle) by leveraging a pre-trained video diffusion model.
@@ 1 Introduction
Humans effortlessly understand how to effectively use objects to achieve their goals within a given environment. This knowledge, known as affordance, originally introduced by Gibson [21], encapsulates a comprehensive understanding of the inherent ways humans interact with objects. Modeling these patterns is crucial for AI systems to assist or mimic human behavior, leading to extensive research in AI fields [93, 85, 79, 92, 57, 72, 40, 1, 63, 61, 64, 78, 65, 15, 10, 121, 9, 113, 123, 97, 53, 62, 71, 122, 12, 75, 36, 55, 11]. However, existing studies on affordance learning primarily focus on static human-object interaction (HOI) patterns. For example, some approaches address affordances by inferring contact (or interaction) region between humans and objects in 2D images [49, 2], 3D humans [90, 110], and 3D objects [109, 110]. Others [26, 37] represent affordances as spatial distributions aggregated from static 3D HOI samples. While recent methods generate human motions including HOI [50, 107, 70, 51, 108, 105, 45] and Human-Scene Interaction (HSI) [100, 112, 105], they often focus on limited actions such as standing, sitting, or picking, without focusing on rich, dynamic nature of object-specific affordances in human motion.
In this paper, we present a novel framework for learning Dynamic Affordance across diverse and unbounded target object categories, without the need for laborious 3D HOI data capture. Our approach is motivated by the observation that human-object interactions involve not only static properties (e.g., contact regions or grasping patterns) but also object-specific characteristic movements over time. For instance, as shown in Fig. 2, humans naturally push a shopping cart and pull a suitcase while walking forward, yet rarely lift their handles vertically. Existing studies, which mainly focus on static affordance features, fail to capture these dynamic interaction patterns. To address this, we present DAViD, a generative motion model for modeling Dynamic Affordance that learns the dynamic and object-specific 4D HOI patterns. To handle the vast diversity of dynamic HOI patterns, our method considers each object-specific 4D HOI pattern as a distinct motion concept, adapting a pre-trained motion diffusion model (MDM) [88] via Low-Rank Adaptation (LoRA) [31]. Once human motions are generated by our MDM with LoRA, we further synthesize object motions conditioned by human motion using a score-based diffusion model, where we introduce a novel sampling technique, named Temporal Guidance Sampling (TGS), for enhancing motion consistency and fine details. Notably, our approach goes beyond modeling dynamic affordance with limited data samples, enables to naturally blend the newly learned HOI motion concepts with motions already known from the pre-trained model in a zero-shot manner, even for multiple HOI concepts.
To train DAViD, we also present a pipeline to synthetically produce realistic 4D HOI samples for a target object category, enabling effective model training without requiring expensive real-world 3D data collection. Most existing dynamic HOI learning approaches [98, 33, 116] rely on the 3D datasets captured in lab-controlled environments [3, 32, 28, 119, 86, 18, 38]. This constraint significantly limits the opportunity to learn diverse and distinctive dynamic affordance patterns across a broader range of object categories. Our key insight is that pre-trained video generation models inherently capture dynamic patterns of object usage. As shown in Fig. 2, these models already know plausible dynamic interactions (e.g., pushing a shopping cart or pulling a suitcase). Since it presents non-trivial challenges to leverage the synthetic 2D videos for learning 4D dynamic affordance, we propose a novel pipeline that synthesizes plausible 4D HOI samples for a given 3D object mesh. We achieve this by generating 2D videos aligned to the 3D object, and aligning the uplifted 4D human motion sequences with the object model, guided by synthetic video cues.
We validate the efficacy of our approach by learning Dynamic Affordance on 30 3D object categories, gathered from multiple sources [3, 32, 8, 106, 60, 37, 83]. We compare our DAViD with other HOI motion synthesis approaches [50, 51, 88] against public dataset [50], demonstrating that our model outperforms competitors in generating HOI motion with plausible contact. We further demonstrate the advantage of our pipeline, by generating results for a various previously unexplored target objects, and also validating its advantage in merging multiple HOI motion concepts with known humn motion from pre-trained MDM [88].
In summary, our main contributions are as follows: (1) We present a generative motion model, DAViD, to learn Dynamic Affordance; (2) We present a pipeline to synthetically generate diverse 4D HOI samples leveraging video diffusion models; (3) We address generating various novel HOI motions by fusing multiple, separately trained LoRAs with known human motions from a pre-trained model. The results and our source code will be released.
@@
Figure 2: Dynamic Patterns in HOI. People show dynamic patterns during HOI. Pre-trained video diffusion models have the knowledge of the dynamic patterns.
@@ 2 Related Work
Reasoning Visual Affordance. Affordance is a concept defined by Gibson [21] as the set of possible interactions that an agent can perform within an environment. As the concept narrows to the primary agent (or humans) and the specific interaction target (or objects), the area of learning affordance mainly focuses on HOI. Specifically, followed by the intuition that the visual geometry of an object is closely related to its functionality, many studies leverage visual cues to reason about object’s affordance. The studies are subdivided into various areas, including affordance classification [93, 85, 79, 92, 57, 72, 40], affordance detection [1, 63, 61, 64, 78, 65, 15, 10], and affordance segmentation [75, 36, 55, 11]. Such traditional studies infer affordance for the given context (mostly provided as images), as a form of action labels [5, 30, 47], and contact (or interaction) regions in 2D [49, 2] or 3D [90, 109]. Later studies model static HOI patterns (e.g., contact) during interactions [13, 26, 37], going beyond the understanding of individual HOI situations. Recent studies synthesize human motion including HOI [107, 70, 100, 112, 51, 108, 105], showing potential to learn dynamic HOI patterns, but they focus on a text-driven approach, generating quite consistent human motions regardless of object categories (e.g., dragging a chair, holding a backpack). In contrast, our method focuses on learning rich, dynamic object-specific HOI patterns, observable in real-world HOIs.
@@
Learning Concepts in Diffusion Models. Concept learning, also known as concept customization, is a strategy for extending the space of the pre-trained model to understand specific concepts (e.g., subjects, styles). Some studies tune the embeddings [20, 95] corresponding to a given concept in a pre-trained diffusion model, while others [76, 43] tune both the model weights and embeddings together. Recently, LoRA [31], originally used in large language models (LLMs), shows impressive results in producing images well aligned to the given concept [77], leading to an increase of attempts [80, 25, 82] to use LoRA for concept learning. Such concepts can be provided in various forms, including images [52, 120, 54, 23], videos [102, 25], and audios [96]. However, concept learning in fields beyond images, video, and audio is relatively less explored. We apply LoRA to the human motion space of pre-trained MDM [88] to learn the dynamic patterns displayed in HOI as a single concept, extending the original human motion space in terms of HOI.
@@ 3 Method
Our approach aims to learn the Dynamic Affordance of a target object from synthetically generated 2D videos, which are produced by a video diffusion model [41]. In the first stage, we generate diverse 4D HOI samples interacting with a given 3D object mesh by leveraging pre-trained video models (Sec. 3.1). Then, our method models the collected 4D HOI samples into our DAViD, which consists of (1) a motion diffusion model along with LoRA, and (2) an object pose diffusion model. The first module synthesizes feasible human motions for the target object, and the second module synthesizes object motions paired with previously generated human motion, using our novel sampling technique, TGS (Sec 3.2). An overview of our method is shown in Fig. 3.
@@
Our pipeline begins with a 3D object mesh input, and produces diverse 2D HOI videos showing how a person interacts with (or uses) the target object.
2D HOI Video Synthesis. Given an input 3D object mesh, we produce multiple 2D HOI videos capturing human-object interactions with the object. We first render the object from cameras with multiple viewpoints. The intrinsic parameter is fixed to match the same perspective camera model and parameters of the world-grounded human mesh recovery method (GVHMR [81]) used in our later 3D lifting process, ensuring a consistent camera setup. For the extrinsic parameters, we consider diverse viewpoints based on the object’s type, either ground-placed object or portable. For relatively stationary object (e.g., cart), we consider horizontal viewpoint variations centered on the object, while for the portable objects (e.g., trumpet) we sample viewpoints within a specified range from all view angles. The object type is automatically inferred via a vision-language model [67], by querying with the rendered image with a prompt: “Is the object plausibly on the ground while being used?”.
Once the object is rendered from each view, we synthesize a realistic human naturally interacting with it using an off-the-shelf pre-trained 2D image diffusion model [44]. Unlike the previous similar approaches [42, 111, 37] that use inpainting by providing a mask region to a model, we use ControlNet [115] to synthesize the entire image, including the human and the plausible background, by using edge cues of rendered object [6] as the condition. Our primary objective is to synthesize both plausible human-object interaction for the given object viewpoint and a realistic background, which is important for estimating dynamic cues and camera poses in a global coordinate. After synthesizing HOI images from several views, we refine the results by filtering out low-quality or undesirable images. We find it is beneficial to use small amount of human effort for this, taking approximately three minutes per category.
Finally, we synthesize corresponding 2D HOI videos ${\{\mathcal{V}_{d}\}_{d=1}^{N}}$ from each 2D HOI images ${\{\mathcal{I}_{d}\}_{d=1}^{N}}$ by using a pre-trained image-to-video generation model [41] with the same text prompt used for generating images [^1] As we synthesize several images at each camera view, the number of videos $N$ is not necessary the same as camera view number, $C$.. Notably, our synthesized 2D HOI videos have key advantages over in-the-wild 2D real videos in learning dynamic affordance: (1) our HOI videos contain diverse views and interaction scenarios for the input 3D object, (2) we have precise camera parameters and 3D object pose information for the initial frame of each video. These are important factors for reliable 4D HOI sample generations.
@@
Uplifting for 4D HOI Sample Generations. For each 2D HOI video $\mathcal{V}$ 222we drop the video index $d$ for simplicity., we uplift 3D human motions and 3D object movements to generate 4D HOI samples. Our key strategy is to uplift each component (human, object) by finding each of the relative camera poses over time, and co-locate them in the common camera coordinate shared by both components.
For the object side, we use Perspective-n-Points (PnP) [19, 48] algorithm to find the relative camera pose with respect to the given 3D object mesh. Specifically, we aim to compute the extrinsic camera poses $\{\bm{R}_{o}^{f},\bm{t}_{o}^{f}\}_{f=1}^{F}$, where the one for the first frame $f=1$ is already known. We first identify visible object mesh vertices $\{\mathbf{V}_{o}\}_{\text{vis}}$ via raycasting [74] for the first frame $f=1$, where $\mathbf{V}_{o}\in\mathbb{R}^{3}$ is an object vertex. The 2D projections of the visible vertices at the first frame is denoted as $\{\bm{v}_{o}^{f=1}\}_{\text{vis}}$, where $\bm{v}_{o}^{f=1}\in\mathbb{R}^{2}$. By applying an off-the-shelf 2D video tracker [35, 34], we obtain the corresponding 2D points $\bm{v}_{o}^{f}$ in the remaining frames $f>1$.
As we have the correspondences between 3D object vertices and matched 2D points for every frame, we apply PnP [19, 48] to compute the camera poses at frame $f$ with respect to the object mesh in its canonical space:
$(\bm{R}_{o}^{f},\bm{t}_{o}^{f})=\text{PnP}\left(\{\bm{v}_{o}^{f}\}_{\text{vis}},\{\bm{V}_{o}\}_{\text{vis}}\right).$ (1)
For the human side, we leverage world-grounded human motion recovery models, GVHMR [81], denoted as $\mathbf{F}_{\text{human}}$ to uplift human motions into 3D from the synthesized 2D HOI videos. We use GVHMR [81], except in the cases where human translation occurs without walking motions (e.g., riding a motorcycle). In such cases, we use TRAM [99], as GVHMR [81] typically fails to predict plausible camera motion under these conditions. Therefore,
$\mathbf{F}_{\text{human}}(\mathcal{V})=\{\bm{\theta}^{f},\bm{\beta}^{f},\bm{\phi}^{f},\bm{\tau}^{f},\bm{R}_{h}^{f},\bm{t}_{h}^{f}\}_{f=1}^{F},$ (2)